Operations Report – November 2nd

SOL: 4
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: No
Currently charging: Yes
Opportunity rover used: Yes
Hours: 143.0
Beginning charge: 100%
Ending charge: 90%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 256.3
Beginning charge: 100%
Ending charge: 69%
Currently charging: Yes
Perseverance rover used: No
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: 19 gallons
Water (static tank): 387 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: An extra window was requested for today as we had an interview on live TV, both audio and video worked smoothly without any issues.
Summary of suits and radios: A single one-piece suit was lent to the filming crew. It was later cleaned and disinfected. No radio was given to them. We wish to note that the actor wearing it bumped it a couple of times while trying to enter the rover for the video. We will inspect it tomorrow in search of any new marks on the acrylic.
Summary of GreenHab operations: measurements and watering
WATER USE: 15 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: same as yesterday: used the blender, 3 large and 1 small paper filters, the microwave, the oven, 2 large erlenmeyer flasks, 1 large and 4 small beaker glasses, the scale, and the mortar and pestle. All this was used to recover starch from rice and potatoes and then mix it with a soil sample to create a concrete-like material. It was also used by the filming crew in the morning.
Dual split: Off
Summary of RAM operations: Used only by the filming crew
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: As reported by the commander over the radio, today after the filming crew left we found toilet paper inside the toilet bowl and wet wipes in the burnable trash bin.

Journalist Report – November 2nd

by Marina Buqueras Rovira


Streaming y Film en Marte

Sol 4

9 am Testeando las comunicaciones con la Tierra, todo funciona bien con la TV BDN!
La Universidad de Purdue llegó y David nuestro Commander los recibió en nuestra casa.

10 am Un buen desayuno de pancakes con arándanos y chocolate nos hace perder la noción.

12 pm La tripulación se pone de gala y ya estamos listos para la entrevista en vivo de la televisión Badalona Comunicació. La buena vibra nos mantiene con humor, el Commander y el XO, hacen que la Crew 285 sea un equipo de diez!

14 pm La sopa caliente nos recompone de la ardua sensación de nostalgia, después de nuestro directo, ya que sabemos que nuestros familiares nos han visto y esperemos que les lleguen nuestros abrazos.

16 pm La tarde, en el SciencieDome, la dedicamos a nuestros proyectos, Andrea sigue con sus mezclas de materiales de Marte para su experimento y Luis preparándose para su próxima observación solar y su astrofotografía del día, Andrés manteniendo y plantando nuevas semillas en el GreenHab, Marina les acompaña para captar todos estos momentos.

Tom, hoy fue el nuevo actor de la MDRS, nuestro Engineer és el nuevo Matt Damon en The Martian!

Mañana nos espera un día intenso con gran actividad, así que acabamos el sol 4 preparando bien los EVA.

Seguimos fuertes!


Streaming and Film on Mars

Sol 4

9 am Testing communications with Earth, everything is working well with TV BDN!
Purdue University has arrived, and our Commander David welcomed them to our home.

10 am A delicious breakfast of blueberry and chocolate pancakes makes us lose track of time.

12 pm The crew dresses up, and we are all set for the live interview on Badalona Comunicació TV. The good vibes keep us in high spirits; the Commander and the XO make Crew 285 a top-notch team!

2 pm Hot soup revitalizes us from the intense feeling of nostalgia after our live session, knowing that our families have watched, and we hope they feel our hugs reaching them.

4 pm The afternoon at the ScienceDome is dedicated to our projects. Andrea continues with her experiments with Martian materials, Luis prepares for his next solar observation and daily astrophotography, Andrés maintains and plants new seeds in the GreenHab, and Marina comes with them to capture all these moments.

Tom, today, took on the role of the new actor at MDRS; our Engineer is the new Matt Damon in "The Martian"!

Tomorrow, we anticipate a busy day with significant activity, so we conclude Sol 4 by preparing well for the EVAs.

We remain strong!

Supplemental Operations Report – November 5th

Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Rovers: Perseverance, Curiosity, Spirit, Sojourner. A tunnel stick on the right from the rear airlock.

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 19 %; Avg 54.4%.
VDC Last 24 hours: Max 58.58V; Min 43.59V; Avg 50.65V.
Generator run time: 3928.8 hours. Oil changed on 11/05/2023

Propane Readings:
Station Tank: 79%
Director Tank: 82%
Intern Tank: 85%
Generator Tank: 80%"

Hab Static Tank – 347 gallons
GreenHab – 168 gallons
Outpost tank – 350 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: No.
Hours: 195.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: All rovers still waiting for all parts to arrive.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Hanksville for supplies.
General notes and comments: All cars awaiting renewed insurance ID.

Summary of Internet: All nominal.
EVA suits and radios:
Suits: All nominal
Comms: All nominal

Campus wide inspection, if action taken, what and why: One tunnel stick on the right from the rear airlock is corroded out at the ground level and needs to be replaced.
Summary of Hab Operations: Toilet tank level sensors need to be cleaned from toilet paper.
Summary of GreenHab Operations: All nominal
Summary of SciDome Operations: All nominal
Summary of Observatories Operations: Robotic observatory had a mount error but currently is operational. Solar observatory but nominal.
Summary of RAM Operations: Remote temperature sensor for smart home systems is offline. Most likely due to the battery being dead and it needs to be relocated to avoid touching metal walls.
Summary of Outpost Operations: Directors trailer thermostat was randomly offline three times. A new one is ordered to diagnose this issue.
Summary of Health and Safety Issues: All nominal.

Copyright © The Mars Society. All rights reserved. | Main Site